OpenStudy (anonymous):

WILL GIVE BEST ANSWER What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -83? -200 -197 -194 -191

OpenStudy (anonymous):

Can you help me with this?

OpenStudy (evonhowell):

Yeah I'm working on it:)

OpenStudy (anonymous):

THANK YOU. XD

OpenStudy (evonhowell):

a1 = a17 - 16d 13 = -83 - 16d d =?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

where does the 16d come from?

OpenStudy (evonhowell):

35th term = a1 + (n-1)d = -6 + (35-1) d = -6 + (34) (?) =

OpenStudy (evonhowell):

I believe that is how it is worked haven't done this in awhile

OpenStudy (anonymous):

-6 + (34) (?) = Do i have to find the ? and the =__ ?

OpenStudy (evonhowell):

Yes I'm not going to just give you the answer LOL

OpenStudy (anonymous):

I know, XD. thanks for the help.

OpenStudy (evonhowell):

Your Welcome:)

OpenStudy (evonhowell):

Hope I helped out a little bit:)

OpenStudy (anonymous):

You definitely did. :D